### v1

# Revisiting Privacy, Utility, and Efficiency Trade-offs when Fine-Tuning Large Language Models

Abstract
We study the inherent trade-offs in minimizing privacy risks and maximizing utility, while maintaining high computational efficiency, when fine-tuning large language models (LLMs). A number of recent works in privacy research have attempted to mitigate privacy risks posed by memorizing fine-tuning data by using differentially private training methods (e.g., DP), albeit at a significantly higher computational cost (inefficiency). In parallel, several works in systems research have focussed on developing (parameter) efficient fine-tuning methods (e.g., LoRA), but few works, if any, investigated whether such efficient methods enhance or diminish privacy risks. In this paper, we investigate this gap and arrive at a surprising conclusion: efficient fine-tuning methods like LoRA mitigate privacy-risks similar to private fine-tuning methods like DP. Our empirical finding directly contradicts prevailing wisdom that privacy and efficiency objectives are at odds during fine-tuning. Our finding is established by (a) carefully defining measures of privacy and utility that distinguish between memorizing sensitive and non-sensitive tokens in training and test datasets used in fine-tuning and (b) extensive evaluations using multiple open-source language models from Pythia, Gemma, and Llama families and different domain-specific datasets.
1 Introduction
Large language models (LLMs) have demonstrated remarkable abilities in solving varied natural language tasks [42], with a host of applications in multiple domains such as education [48], medical chatbot [46], AI assistant [10], etc. Their ability is acquired by first pre-training models on a large Web corpus[23], which results in general language understanding [6] and then fine-tuning the pre-trained models on domain-specifc data, which customizes their performance to specific tasks [55]. Two foundational challenges with pre-training/fine-tuning these models at scale are centered around enhancing privacy and efficiency, while retaining the models’ utility. The former is related to reducing the risk of LLMs leaking sensitive user information contained in the training data, while the latter is related to reducing the computational cost of training.
A long line of recent research in the privacy community has focussed on methods to mitigate privacy risks when training LLMs [44, 56, 1, 51]. A notable example of such methods is differentially privacy (DP)[1]. Similarly, a flurry of recent research in the systems community has explored parameter-efficient fine-tuning (PEFT) methods in LLMs [14]. A notable example of this class of methods is low-rank adaptation (LoRA) [20]. However, no prior works, to the best of our knowledge, have investigated the privacy risks associated with the efficient training methods. So the central question driving our research here is: do efficient fine-tuning methods enhance or mitigate privacy risks during training?
Our question above is motivated by the following high-level observations about DP and LoRA: methodologically, both DP and LoRA restrict the impact training examples can have on model parameters – DP deliberately through its noisy gradient update, and LoRA through low-rank adaptation. However, DP incurs significant additional computational overhead [11] compared to traditional fine-tuning, while LoRA significantly reduces the computational costs. So the answer to the above question can have significant consequences for achieving good privacy-efficiency-utility tradeoffs when fine-tuning. For instance, if LoRA mitigates privacy risks of training, then it suggests that this method can simultaneously achieve both privacy and efficiency objectives, contradicting the conventional wisdom drawn from DP literature that privacy comes at a computational cost. One of the key (surprising) findings of our work below lies in establishing that LoRA does indeed mitigate privacy risks.
When attempting to answer the above question, we encountered a more foundational question that has been not been well-addressed by the existing literature namely, how should one quantify the privacy risks associated with a fine-tuning method, so that it allows for a performance comparison across different methods? For instance, the claimed privacy benefits of the DP method are based on theoretical guarantees and not empirical evaluation, making it hard to compare it against privacy benefits of other fine-tuning methods.
Numerous recent works explored privacy concerns with LLMs that arise from their ability to memorize and recollect training data sequences containing sensitive personally identifiable information (PII), such as names, email addresses, phone numbers, credentials, etc. This sensitive information can leak from an LLM, either by accidental generation in response to benign requests to the model, or by deliberate extraction using adversarially crafted requests to the model[40, 38, 12, 22, 9, 41, 43].
However, surprisingly, existing privacy as well as utility metrics for LLMs ignore the distinction between sensitive and non-sensitive counterparts (tokens) in training data sequences [7, 40, 3]. [^1] A few works used this distinction to train models [44, 56], but it has not been applied when quantifying privacy risks or utility metrics. Sensitive data is often inherently random and less predictable, such as social security number or phone number. Non-sensitive data, on the other hand, is more predictable, e.g. there is only a small set of valid completions for the sentence “The dog chases the _". Consequently, the ability of LLMs to recollect or predict sensitive vs. non-sensitive data is very different (see Figure 1 for details). As we show in Section 3.2, ignoring this distinction leads to inaccurate measures of privacy and utility, and in particular, may have resulted in some prior works unintentionally exaggerating the threat of LLM’s memorizing training data[3, 9, 7].
Here, we propose and justify new measures for privacy and utility explicitly accounting for the differences in model’s performance for sensitive and non-sensitive data. For sensitive data, we are primarily interested in privacy, i.e. we want the loss on sensitive tokens to be as high as possible. Conversely, for non-sensitive data, we are interested in utility, i.e. we want the loss on non-sensitive tokens to be as low as possible.
(a) Pretrained
(b) Fine-tuned
Figure 1: Sensitive and non-sensitive tokens have different predictability, measured as the recollection loss by pre-trained models in Figure 1(a) and fine-tuned models (recorded at epoch $5$) in Figure 1(b). The distinction motivates us to quantify privacy using sensitive token loss on training data (higher is better) and utility as non-sensitive token loss on test data (lower is better).
Contributions. We summarize our main findings below:
1. Quantifying privacy and utility, when training LLMs: We conceptually argue and empirically demonstrate that LLM’s ability to recollect (predict) sensitive and non-sensitive data in training (test) datasets are so starkly different that we need to account for them when quantifying privacy and utility. Privacy is best captured by a model’s ability to recollect sensitive tokens in training data, while utility is best captured by the model’s ability to predict non-sensitive tokens in test data.
2. Exaggerated estimates of privacy threat from LLM memorisation: We show that compared to our measure, existing measures significantly over-estimate privacy risks from LLM training. LLMs recollect non-sensitive tokens at a significantly higher level than sensitive tokens. So by not distinguishing between them, existing studies are over-estimating privacy threats.
3. Comparing privacy-utility-efficiency tradeoffs for three different fine-tuning methods: Our measures allow us to conduct a systematic and extensive empirical study of three different fine-tuning methods: full fine-tuning, DP, and LoRA, using models from three different LLM families, Pythia [4], Gemma [45] and Llama [47] over both real-world and synthetically-generated datasets. For each method and model, we evaluate privacy, utility (in terms of train and test loss respectively) and its impact on benchmark performance, and efficiency. Our comparative study yields several interesting insights: we find that – full fine-tuning results in poor utility-privacy tradeoffs; DP offers the best utility-privacy tradeoffs, but is computationally very expensive; LoRA is almost on par with DP in terms of utility and privacy, but is much more computationally efficient.
4. Feasibility of achieving all the three privacy, utility and efficiency objectives simultaneously: Our findings about LoRA performance suggest that it may be possible for methods to achieve all three objectives simultaneously. Our results challenge prevailing wisdom that enhancing privacy during training is necessarily more computationally expensive and call for investigating privacy benefits of existing and new parameter efficient fine-tuning methods.
2 Related Work
Privacy Attacks and Quantification. Privacy concerns in LLMs have gained attention in recent years [24, 49], particularly the possibility of exposing data via membership inference attacks (MIA) [40, 38, 12, 22] and training data extraction attacks [9, 41, 43]. While membership inference attacks have the goal of identifying if a certain datapoint was in the training dataset using model confidence scores, training data extraction attacks try to extract specific parts of data from the training dataset using different prompting strategies. A recent study [54] underscores that membership inference attacks may not be reliable due to their practical limitations of requiring knowledge of the entire training data, thus proving data extraction attacks to be more viable.
Existing works have quantified privacy attacks in large language models especially through the lens of memorization. For example, the authors in [7] assess memorization by quantifying how closely LLM generation matches the exact training data phrases when prompted with tailored prefixes. The metric of ‘exposure’ is introduced in [8] to evaluate a model’s vulnerability when exposed to data that was artificially introduced into the dataset (also known as "canaries") several times during training. This metric has been widely adopted in subsequent research [44, 56] as a measure of privacy. However, the metric’s formulation is limited by its reliance on assumptions about the surrounding knowledge of other canaries, which do not hold in practical scenarios.
Intuitively, privacy evaluation should be more related to sensitive data seen by the model during its training. Some of the existing studies  [25, 36, 56, 44] focusing on privacy attacks in LLMs may consider sensitive and non-sensitive counterparts in the data explicitly during training, but they do not retain this distinction during privacy quantification. We believe (and later show) that this lack of distinction can lead to an inaccurate (sometimes, overestimated) assessment of privacy threats. In our paper, we carefully make the distinction between sensitive and non-sensitive data and propose a revised quantification.
Privacy-Utility Tradeoffs. To mitigate privacy leakage, differential privacy (DP) measures have been proposed, which add theoretical privacy guarantees to the training process [1]. The key idea of DP is to clip the gradients of each datapoint and add noise to the clipped gradients in every iteration. Its primary goal is to reduce the influence of individual datapoints in the training procedure thus preventing its leakage during inference.
Existing work utilising DP [51] has shown a clear trade-off between privacy and utility where privacy is measured in terms of theoretical guarantee and utility in terms of overall performance on the end task. The authors in [5] also question the notion of privacy considered in these privacy preserving techniques. Another line of work from [44] and [56] distinguishes between sensitive and non-sensitive data using techniques like regex and redaction and proposes a customised training strategy using DP. However, their evaluation measure also relies on theoretical guarantees for privacy and overall performance for utility. The question that arises here is whether this measure of quantification of privacy and utility is sufficient in the context of large language models, or whether it needs to be more nuanced.
Additionally, to ensure privacy guarantees, differential privacy requires higher computational resources (time and memory) during training. For example, compared to vanilla SGD, DP-SGD may incur up to $20$x training time, which is often a bottleneck in resource-intensive tasks [11]. Therefore, the conventional wisdom is that privacy comes at the cost of efficiency. This leads us to probe ways in which the above limitation can be prevented.
Utility-Efficiency Tradeoffs. Fine-tuning a pre-trained LLM on a task specific context has been used in several applications like medical chatbots, AI assistants, etc. However, full fine-tuning of large language models is expensive as it involves updating all the parameters. Recent work has focused on reducing the training cost while maintaining utility, with the introduction of parameter efficient fine-tuning (PEFT) techniques [14]. Well-known PEFT methods include adapter-based fine-tuning [19, 27, 57, 15], soft prompt-based fine-tuning [30, 29, 33, 34, 28], and parameterized fine-tuning [20, 32]. Among different PEFT methods, the Low-Rank Adaptation method, called LoRA [20] has emerged as one of the most widely used methods. LoRA updates fewer parameters in the model via low-rank approximation, providing computational efficiency with a relatively low cost to utility.
In recent work on private fine-tuning [35, 50, 37], the authors combine LoRA with DP to reduce the additional computational overhead induced by differential privacy. In such a context, we ask whether DP is the only method towards ensuring privacy. LoRA’s training procedure of updating fewer parameters can be thought of as analogous to the noisy gradient update in DP. This leads us to the question of whether LoRA has any privacy benefits besides having control over utility-efficiency tradeoffs.
Privacy-Utility-Efficiency Tradeoffs. To the best of our knowledge, our work is the first to investigate the privacy benefits of LoRA and systematically examine the privacy, utility, and efficiency tradeoffs among different fine-tuning methods. Besides, instead of relying on the existing measures of privacy and utility, we distinguish between sensitive and non-sensitive data to propose a nuanced quantification of privacy and utility.
3 Quantifying Privacy and Utility
In this section, we introduce the distinction between sensitive and non-sensitive entities when quantifying privacy and utility of an LLM. We conduct case studies to compare our quantification with existing measures. Finally, we demonstrate how the privacy threat is unintentionally exaggerated in existing methods due to the lack of distinction between sensitive and non-sensitive entities.
3.1 Rethinking Privacy and Utility
Existing studies at the intersection of privacy and natural language processing [56, 44, 31, 52] seek to enhance privacy while maintaining model utility. Utility is generally assessed based on model performance, such as loss, accuracy, or perplexity across the entire test dataset. Privacy is evaluated in terms of performance measures on the entire training dataset or theoretical guarantees in differential privacy (DP).
Natural language text may contain both sensitive and non-sensitive words, referred to as entities. For example, sensitive entities include names, addresses, phone numbers, order IDs, and other personally identifiable information. In contrast, non-sensitive entities generally involve semantic and/or syntactic completions following predictable patterns in language generation tasks. Informally, sensitive entities are drawn from a large search space (e.g., a random sequence of digits), resulting in high entropy and low predictability. In contrast, non-sensitive entities are more restricted in their occurrences (e.g., a subject is typically followed by a verb), leading to low entropy and high predictability. Several studies [3, 44, 56] distinguish between sensitive and non-sensitive entities in their proposed privacy leakage mitigation methods. However, the distinction is not leveraged in the quantification of privacy and utility, which is essential for a granular evaluation as discussed next.
Quantification of privacy and utility. In this work, we quantify privacy and utility by accounting for sensitive and non-sensitive entities. Considering a training dataset and a test dataset in a general LLM training pipeline, we quantify privacy as the recollection of sensitive entities in the training data and utility as the prediction of non-sensitive entities in the test data. Our motivation for the quantification is two-fold: (1) privacy of a model is generally related to training data, while utility is the model’s performance on the test data. (2) when quantifying privacy, we deliberately disregard non-sensitive entities, since they are more predictable and not sensitive to a specific person or entity. Similarly, in quantifying utility, we ignore sensitive entities in the test data, since the sensitive entities are rare (and possibly unseen during training), whereas predicting non-sensitive entities indicates the general language understanding ability of LLMs. Next, we provide two pieces of evidence supporting why the distinction is important.
3.2 Why do we distinguish between sensitive and non-sensitive entities?
In this section, we present evidence supporting the importance of distinguishing between sensitive and non-sensitive entities in natural language text while quantifying privacy and utility of an LLM.
(a) Privacy measure in Pythia
(b) Utility measure in Pythia
(c) Privacy measure in Gemma
(d) Utility measure in Gemma
(e) Privacy measure in Llama2
(f) Utility measure in Llama2
Figure 2: Our measures offer a more precise assessment of privacy and utility when fine-tuning LLMs by distinguishing between sensitive and non-sensitive tokens, revealing higher privacy (higher loss) for sensitive tokens and better utility (lower loss) for non-sensitive tokens compared to traditional measures that overlook this sensitivity-based distinction.
I. Analyzing privacy and utility while fine-tuning an LLM
In order to align with LLM terminology, hereafter, we use tokens to denote entities. Fine-tuning involves iterating an LLM on a specific dataset containing both sensitive and non-sensitive tokens. We illustrate how our measure of privacy and utility compares to existing measure in a typical fine-tuning scenario, highlighting a key difference: our approach distinguishes between sensitive and non-sensitive tokens, whereas the existing measure does not.
Results. In Figure 2, we demonstrate measures of privacy (left column) and utility (right column) while fine-tuning three LLM models on Customersim dataset [44] (experimental details are provided at the end of this section). In particular, we show training loss on the left column and test loss on the right column. Importantly, we separately compute the loss for both sensitive and non-sensitive tokens in both training and test datasets. Intuitively, a higher loss denotes more privacy and less utility.
Privacy is overestimated in the existing measure. In Figure 2(a), we compute privacy using our measure, as well as the existing one. The existing measure of privacy considers all tokens in the training data, where low training loss denotes less privacy, while our measure considers only the sensitive tokens in the training data. Using our measure, a notable disparity emerges: sensitive tokens exhibit significantly higher loss than non-sensitive ones, particularly in the initial training epochs, as sensitive tokens are less predictable. This eventually indicates that the loss over all tokens (existing measure) would be much lower initially than the loss over only sensitive tokens (our measure), thus overestimating privacy threats much earlier. Similar trends are observed for other models in Figures 2(c) and 2(e).
Utility is underestimated in the existing measure. In Figure 2(b), we compute utility using our measure and the prevailing one. The existing utility measure is related to the test loss of all tokens, where lower test loss indicates better utility. We can observe that our measure that considers the test loss on only non-sensitive tokens provides better utility than the existing measure. Similar trends are observed for other models in Figures 2(d) and 2(f). Existing measure would lead to better utility at earlier epochs compared to our measure i.e. one would select checkpoints at epochs $4$, $2$ and $2$ using existing measures for Pythia, Gemma and Llama2 respectively and at epochs $5$, $6$, and $6$ using our measure. Thus, utility is underestimated in the existing measure. Figure 18 in Appendix 14 shows the change more distinctly in log-scale. For the same fine-tuning scenario, our measure certifies better utility than the existing measure, which unintentionally considers the loss of sensitive tokens that are rare. The key to our findings is how we annotate sensitive and non-sensitive tokens in real-world datasets. Next, we expand our experimental details leading to our results.
Table 1: Example of samples from datasets with sensitive tokens highlighted.
CustomerSim
SYS: Hello, I am the customer
support bot. What can
I do for you?
USR: Hello robot. I ordered a pot
several days ago but I can’t track it.
SYS: Could you verify your full name?
USR: Patrick Schug
SYS: Verify your order number please.
USR: It’s 843-58572-7002.
SYS: You can track your package
with your tracking number
, which is AGZIM5T6KL.
Are you happy about my answer?
USR: All good. See you.
SYS: Have a nice day! Bye.
SYS: Hello, I am the customer
support bot. What can
I do for you?
USR: Hi. Where is my package?
SYS: Could you verify
your full name?
USR: I am Catherine Pena.
SYS: Verify your phone number.
USR: You can reach me at
547.302.3744.
SYS: The tracking number is
VVTPHDB6VK.
Anything else?
USR: All good.
SynBio
My name is Alexander Tanaka, and I’m a
saleswoman with a year of experience. I
recently completed a project that
involved developing and implementing
a new sales strategy for
my company. I started by analyzing our
current sales data to identify areas where
we could improve…
My name is Phillip Martinez, and
I would like to share some aspects
of my life’s journey with
you. I have had the pleasure of
living in various places throughout
my life, but I currently reside
at 4537 Tanglewood Trail
… you can reach me via email at
phillip-martinez@outlook.com
or by phone at +86 19144 1648.
Experimental setup and methodology. We perform our analysis on two datasets: CustomerSim [44], a simulated dialog dataset for conversation generation and SynBio (originally called PII) [18], an LLM generated dataset representing student biographies containing personal identifiable information. Table 1 shows some excerpts from the datasets. We use three open-source models during evaluation: Pythia 1B [4], Gemma 2B[45], and Llama2 7B [47].
We leverage two tools for annotating sensitive information in a given text: Presidio [39], which helps in identification of private entities in text, and GPT-4 [2], which is provided with a particular prompt for returning the annotated portions. An example of such a prompt for annotating samples is provided in Appendix 11.
We run two surveys, each among 40 Prolific [^2] https://www.prolific.com users, to gauge the usefulness of the two tools. We provide the details of the survey in Appendix 13. Figure 3 shows the results on the CustomerSim dataset which depicts that 75% participants found GPT4’s annotations to be accurate while Presidio annotations were mostly mixed or under-annotated. Hence, throughout the rest of the paper we show our results using GPT-4 annotations, and those using Presidio annotations are shown in Appendix 15.
To summarize, the degree of difference in computing privacy and utility using our measure and the existing one depends on the ratio of sensitive to non-sensitive tokens. A higher ratio would result in a higher difference in the measure, and vice versa. Considering the distinction between sensitive and non-sensitive tokens, we show that the existing measure can both exaggerate privacy threats and underestimate the utility in LLMs. In this context, we re-examine a prior study [3] to better support our claim that reported privacy threats are exaggerated.
Figure 3: GPT-4 shows higher annotation accuracy, with 75% of participants rating its annotations to be accurate while Presidio annotations were mostly mixed or under-annotated.
Table 2: Examples of memorized sequences from [3], often containing predictable and non-sensitive patterns, like mathematical series and licensing text.
Prompt Generation
264. 265. 266. 267. 268. 269. 270. 271. 272. 273. 274. 275. 276. 277. 278. 279.
active.disabled:focus, .datepicker table tr td.active.disabled:hover:focus, .datepicker table tr td.active:active, .datepicker table tr td.active:hover:active, .datepicker table tr td.active.disabled:active, .datepicker table tr td
[…]
amp amp amp amp amp amp amp amp amp amp amp amp amp amp amp amp amp amp amp amp amp amp amp amp amp amp amp amp amp amp amp amp
.word 0 .word 0 .word 0 .word 0 .word 0 .word 0 .word 0 .word 0 .word 0 .word 0 .word 0 .word 0 .word
II. Examining memorized sequences from [3]:
We consider a case study to analyze the reported memorized strings by [3], Our goal is to examine whether the memorized strings contain sensitive information or mere syntactic and semantic patterns.
Experimental setup. The authors in [3] considered the task of predicting whether a model memorizes specific training data points from the Pile dataset [13], which is used to train base LLM models. Among published memorized strings, we randomly choose $5,000$ strings from the pythia-1b-dup split [3]. A representative list of memorized strings is in Table 2, where the strings often follow syntactic and semantic patterns, such as completion of mathematical series, code snippets, licensing agreements, etc. Therefore, our hypothesis is that most of the memorized strings contain a great amount of non-sensitive and highly predictable tokens. To validate our hypothesis, we query for the source of memorized strings with respect to the training dataset, Pile, which aggregates data from multiple sources such as Pile-Cc, OpenWebText, ArXiv, etc. We leverage GPT-4 model to accomplish our task – given a memorized string, we ask for the source of the string from the list of Pile sections. The prompt template for GPT-4 is the following.
{mdframed}
[backgroundcolor=lavender, linewidth=0pt] You are provided with the following text: {memorized–sequence}.
Which section of the Pile dataset does the text belong to? Choose from the list below. You can select 1 or 2 options separated by a comma. Please respond with only the option number.
a. Pile-CC b. PubMed Central c. Books3
d. OpenWebText2 e. ArXiv f. GitHub g. FreeLaw
h. Stack Exchange i. USPTO Backgrounds
j. PubMed Abstracts k. Gutenberg (PG-19)
l. OpenSubtitles m. Wikipedia (en)
n. DM Mathematics o. Ubuntu IRC
p. BookCorpus2 q. EuroParl r. HackerNews
s. YoutubeSubtitles t. PhilPapers
u. NIH ExPorter v. Enron Emails
Figure 4: Memorized sequences are predominantly sourced from GitHub and ArXiv, despite these sections being mid-range in the original Pile dataset, suggesting that memorized content is largely non-sensitive and may pose a lower privacy risk than previously assumed.
Results. In Figure 4, we present two pie charts illustrating the distributions across $22$ distinct sections or data sources within the Pile dataset. The left chart represents the original content distribution of sections within the Pile dataset, while the right chart depicts the distribution of sources of memorized sequences as predicted by GPT-4.
In the right chart in Figure 4, the memorized strings are predicted mostly from GitHub, followed by ArXiv, while the rest of the sources are largely under-represented. Herein, both GitHub and ArXiv are relatively in the middle range in terms of contents in the original dataset in the top chart. However, analyzing the typical data in these sections, GitHub appears as a source of structured format code with repeated predictable patterns, which is commonly tagged as non-sensitive data. Similarly, the Pile dataset includes files uploaded to ArXiv, since is a common typesetting language for scientific research papers [13]. As such, highly memorized strings in the Pile dataset are non-sensitive in nature.
Validating GPT-4 predictions. GPT-4 predictions may be erroneous. Hence, we conduct a verification test to evaluate the accuracy of GPT-4’s predictions. For this assessment, we sample $200$ random strings from each of the 22 sections of the Pile dataset [13], and prompt GPT-4 to predict the source of the strings. Unlike the previous experiment, the ground-truth of string source is known in this validation experiment. Figure 5 illustrates the accuracy for each section, indicating that $50\%$ of the sections exhibit an accuracy rate of at least $90\%$ with $4.5\%$ being the base accuracy of a random predictor. Furthermore, GPT-4 predicts the correct source on an average of $78\%$ strings across all $22$ sections of the Pile dataset. In addition, misclassified strings are often assigned to sections of a similar category, e.g., NIH Explorer misclassified as PubMed Central (more details in Appendix 10). Therefore, the GPT-4 predictions can be considered as reliable.
Figure 5: GPT-4 achieves an average accuracy of 78% in predicting the source of memorized strings across Pile dataset sections, reinforcing the reliability of GPT-4 and supporting our position that privacy concerns in prior work are overestimated without distinguishing token sensitivity.
Figure 6: Most participants classified the memorized sequences detected by [3] as non-sensitive, with fewer than 10% marking them as privacy-sensitive, indicating that the perceived privacy risk of these strings is generally low.
Finally, we conduct a human survey on Prolific to evaluate the extent of sensitive information present in randomly chosen $100$ memorized strings from [3]. The survey results are summarized in Figure 6. The majority of participants classified memorized strings as non-sensitive, while $<10\%$ participants disagree and mark the strings as containing privacy-sensitive information. Therefore, most crowdsourced participants do not perceive the sampled stings as containing privacy-sensitive content. Further details on the setup and results are provided in Appendix 12. Thus, by distinguishing between sensitive and non-sensitive entities, we demonstrate a deeper understanding of actual privacy threat.
4 Privacy-Utility-Efficiency tradeoffs in different fine-tuning methods
(a) CustomerSim
(b) SynBio
Figure 7: Privacy-utility trade-off shows that privacy increases with higher training loss on sensitive tokens, while utility improves with lower test loss on non-sensitive tokens, enabling desired checkpoint selection to balance both objectives.
(a) SCIQ Benchmark
(b) MMLU Benchmark
(c) HellaSwag Benchmark
Figure 8: Full fine-tuning of the Gemma model leads to a significant drop in accuracy on benchmark datasets, with declines of approximately 75%, 9%, and 30% for SCIQ, MMLU, and HellaSwag, respectively.
We use the distinction between sensitive and non-sensitive tokens to study the privacy and utility impact of training models with three different fine-tuning methods: full fine-tuning (FFT), Differentially Privacy (DP), and Low-Rank Adaptation (LoRA). We also investigate the computational efficiency of each method. Our goal is to answer the following questions: “How prone is each method to recollecting the sensitive parts of the training data? (Privacy)” , “How effective is each method at predicting non-sensitive parts of test data? (Utility)” , “What is the computational cost associated with each method? (Efficiency)”.
To answer these questions, we use each fine-tuning method to train three models from different families, Pythia-1B [4], Gemma-2B [45] and Llama2-7B [47], on two datasets, CustomerSim and SynBio. More information about the datasets and our methodology for distinguishing between sensitive and non-sensitive tokens can be found in Section 3.2. We train the models for 50 epochs on each dataset. Details on hyperparameters can be found in Appendix 9.
For each fine-tuning method, model and dataset we report three metrics: 1) privacy as the loss on sensitive tokens (annotated by GPT-4) in the training data, 2) utility as the loss on non-sensitive tokens (the remaining tokens) on a held-out test set, and 3) efficiency based on the relative amount of computation and memory usage of each method. Additionally, to assess how fine-tuning affects the underlying abilities and knowledge of the base model, we measure the performance of the fine-tuned Gemma model (trained with CustomerSim data) on general language understanding benchmarks: SCIQ [21], a dataset of over 13,000 crowdsourced questions on Physics, Chemistry, and Biology; MMLU [17, 16], a large multi-task dataset covering various domains of knowledge; and Hellaswag [53], a dataset for commonsense inference.
Update rules: For each fine-tuning method, we describe how updated weights $W_{t+1}$ are computed from the previous weights $W_{t}$ in each step, where $W_{0}$ are the weights of the pre-trained base model before fine-tuning. We use $X$ to refer to a batch of $|X|$ datapoints and $x_{i}$ to refer to individual datapoint, $\mathcal{M}_{W}$ to refer to the model parameterized by weights $W$, and $\mathcal{L}(\mathcal{M}_{W}(X),X)$ to refer to the autoregressive cross-entropy loss of the model on data $X$. We denote by $\nabla_{W}\mathcal{L}(...)$ as the gradient of the loss wrt. weights $W$ and $\eta$ is the learning rate.
Efficiency: The efficiency of each method is determined by the amount of computation it requires, and also other factors such as memory requirements, which can affect the usable batch-size and thus the overall training throughput. Following [23], we estimate the amount of training compute ($C$) in floating point operations (FLOPs) for full fine-tuning as $C_{\text{FFT}}=6DN$, where $D$ is the number of training tokens and $N$, the number of model parameters. For each method, we report its compute requirements relative to the FFT-baseline based on measurements using the PyTorch profiler [^3] https://pytorch.org/docs/stable/profiler. We also comment on other factors that affect training throughput.
4.1 Method 1 : Full fine-tuning
Update rules: Full fine-tuning (FFT) updates all model parameters at each step:
$W_{t+1}=W_{t}-\eta\nabla_{W_{t}}\mathcal{L}(\mathcal{M}_{W_{t}}(X),X)$ (1)
Privacy-Utility trade-off: Figures 7(a) and 7(b) show the privacy-utility trade-off for the CustomerSim and SynBio datasets, respectively. In these figures, privacy increases with the training loss on sensitive tokens (up $\Uparrow$ on the y-axis), while utility increases when the test loss on non-sensitive tokens decreases (right $\Longrightarrow$ on the x-axis). Each curve starts with the baseline performance of the pre-trained model. For CustomerSim (Figure 7(a)), as training advances (denoted by an arrow $\rightarrow$ on the lines), privacy progressively decreases (lower on the y-axis), while utility improves (rightward on the x-axis) for approximately the first 5 epochs across all models before stabilizing and eventually declining (leftward on the x-axis). However, for SynBio (Figure 7(b)), the privacy-utility trade-off primarily worsens for Gemma and Llama models. On examining these curves, one can select a desired checkpoint that aligns with specified privacy and utility thresholds.
Impact on benchmark datasets: Figures 8(a), 8(b), and 8(c) show the fully fine-tuned Gemma model’s accuracy at each epoch for the three benchmarks: SCIQ, MMLU, and Hellaswag, respectively. Note that the accuracy corresponding to the first point represents the performance of the pre-trained model. We observe that full fine-tuning shows a substantial decline in accuracy (around $0.75$, $0.09$, and $0.3$ decrease in accuracy in SCIQ, MMLU, and HellaSwag, respectively).
Efficiency: FFT serves as our efficiency baseline. It has moderate compute requirements (discussed above), and relatively high memory requirements, since in addition to the input-dependent activations, we need to keep four numbers per model parameter in GPU memory: the parameter value, its gradient, and two optimizer states (first and second moments of the gradient for Adam [26]).
Takeaway: FFT offers poor privacy-utility trade-offs, since gains in utility in most cases come at the cost of a significant loss in privacy. During FFT, models learn to both predict the training distribution better, but also quickly learn to recollect sensitive tokens. In addition, FFT deteriorates the base performance of the model, as can be seen by the rapid decline of the benchmark scores. FFT is moderately efficient and has relatively high memory requirements. The degree of measures along the trade-off, knowledge retention, and efficiency are:
Utility-privacy trade-offs: poor
Retention of base performance: poor
Efficiency: moderate
(a) Pythia (CustomerSim)
(b) Gemma (CustomerSim)
(c) Llama2 (CustomerSim)
(d) Pythia (SynBio)
(e) Gemma (SynBio)
(f) Llama2 (SynBio)
Figure 9: Privacy-utility trade-offs for DP across models and datasets. Top row: For the CustomerSim dataset, DP enhances privacy across all models, though higher noise scales reduce utility, especially for larger models. Bottom row: For the SynBio dataset, DP similarly improves privacy, with utility declines at higher noise levels. Gemma shows a unique utility drop after two epochs, highlighting the complexity of privacy-utility trade-off.
4.2 Method 2 : Differential Privacy
Differential Privacy (DP) algorithms  [1] aim to safeguard the privacy of individual training data points by limiting their influence on the gradient updates during training.
Update rules: DP clips the $l_{2}$ norm of each data point’s gradient at a threshold $T$, followed by adding noise with magnitude $\sigma$ to each clipped gradient. The purpose of clipping is to reduce data sensitivity by ensuring that the impact of data points with high gradient magnitudes on the model parameters is limited. Adding noise further obscures the contribution of individual data points, making it difficult to infer specific data points from the model.
$\begin{split}W_{t+1}&=W_{t}-\eta\,\text{Noise}\left(\frac{1}{B}\sum_{i}\text{Clip}\left(\nabla_{W_{t}}\mathcal{L}(\mathcal{M}_{W_{t}}(x_{i}),x_{i})\right)\right)\\ \text{Clip}(y)&=y/\max\left(1,\frac{\lVert y\rVert_{2}}{T}\right)\\ \text{Noise}(y)&=y+\mathcal{N}(0,\sigma^{2}T^{2}\mathds{1})\end{split}$ (2)
We vary the noise hyperparameter $\sigma$ for the experiments. The clipping gradient norm $T$ is fixed at $10^{-2}$, as in [44].
(a) SCIQ Benchmark
(b) MMLU Benchmark
(c) HellaSwag Benchmark
Figure 10: Fine-tuning Gemma with DP results in a substantial accuracy declines of 70%, 10%, and 30% across the benchmarks.
Privacy-Utility trade-off: Figures 9(a), 9(b), and 9(c) illustrate the privacy-utility trade-off when varying the noise $\sigma$ on the CustomerSim dataset across the Pythia, Gemma, and Llama2 models. As in previous plots, each curve begins with the performance of the pre-trained model. It is evident that DP maintains privacy effectively with minimal degradation. But there is a trade-off between privacy and utility. For all models, lower noise values such as $\sigma=0.1$ are able to achieve better utility than higher ones such as $\sigma=0.5$ and $\sigma=0.9$, but they also decrease privacy more. While the total amount of utility achievable with DP is limited, especially for the larger Llama2-7B model, overall, it provides good privacy-utility trade-offs. Similar trends can be observed for the SynBio dataset in Figures 9(d) and 9(f), but for Figure 9(e) (with Gemma) utility declines after approximately two epochs.
Impact on benchmark datasets: Figures 10(a), 10(b) and 10(c) show the benchmark accuracy over epochs for the DP fine-tuned Gemma model with $\sigma=0.1$ on SCIQ, MMLU, and HellaSwag datasets, respectively. Accuracy drops gradually with increasing epochs for all benchmarks, stabilizing at lower levels, indicating that fine-tuning with DP significantly reduces the knowledge retention capacity.
Efficiency: Differential privacy comes with a high computational cost, since the gradients of each datapoint need norm and clipping computations, and additional noise values are added. Empirically, we observe a relative FLOPs requirement of $C_{\text{DP}}/C_{\text{FFT}}=1.33$. Additionally, the per-sample operations required for clipping mean that we need to keep copies of the gradient for each datapoint in memory, which requires substantially more GPU memory, which decreases the feasible batch size and the overall training throughput even further.
(a) Pythia (Rank 16)
(b) Gemma (Rank 16)
(c) Llama2 (Rank 16)
(d) Pythia (Rank 32)
(e) Gemma (Rank 32)
(f) Llama2 (Rank 32)
Figure 11: Privacy-utility trade-offs for LoRA fine-tuning on the CustomerSim dataset. Top row: LoRA with rank 16 shows that smaller models achieve a better privacy-utility trade-off, while larger models retain utility but experience reduced privacy with more epochs. Bottom row: LoRA with rank 32 yields similar results, with smaller models performing better in privacy-utility trade-off and larger models maintaining utility but with privacy reduction as epochs increase.
Takeaway: DP offers a reasonable privacy-utility trade-off and is less susceptible to learning sensitive data. Increasing noise rates lead to poor utility across all models. Convergence with DP is not guaranteed, particularly in larger models such as Llama2. The privacy gains come at the cost of efficiency. Additionally, fine-tuning with DP leads to quick decline in the benchmark performance. The degree of measures along the tradeoff, knowledge retention, and efficiency are:
Utility-privacy trade-offs: moderate
Retention of base performance: poor
Efficiency: poor
(a) Pythia (Rank 16)
(b) Gemma (Rank 16)
(c) Llama2 (Rank 16)
(d) Pythia (Rank 32)
(e) Gemma (Rank 32)
(f) Llama2 (Rank 32)
Figure 12: Privacy-utility trade-offs for LoRA fine-tuning on the SynBio dataset. Top row: LoRA with rank 16 shows that smaller models achieve a better privacy-utility trade-off, while larger models retain utility but experience reduced privacy with more epochs. Bottom row: LoRA with rank 32 yields similar results, with smaller models performing better in privacy-utility trade-off and larger models maintaining utility but with privacy reduction as epochs increase.
(a) SCIQ Benchmark
(b) MMLU Benchmark
(c) HellaSwag Benchmark
Figure 13: LoRA fine-tuned model maintains accuracy levels close to the pre-trained model with declines of 5%, 3% and 3% across SCIQ, MMLU, and Hellaswag benchmarks, highlighting its effectiveness in knowledge retention.
4.3 Method 3 : Low-Rank Adaptation (LoRA)
LoRA [20] is a parameter-efficient fine-tuning method developed to reduce the compute and memory requirements of fine-tuning LLMs, and to reduce the size of storing fine-tuned checkpoints. It enjoys large popularity for LLM fine-tuning.
Update rules:
$\begin{split}W_{t+1}&=W_{0}+\frac{\alpha}{r}\Delta W_{t+1}\\ \Delta W_{t+1}&=\Delta W_{t}-\eta\nabla_{\Delta W_{t}}\mathcal{L}(\mathcal{M}_{W_{t}}(X),X)\end{split}$ (3)
LoRA freezes the weights of the pretrained base model $W_{0}$ and only fine-tunes an adapter matrix $\Delta W_{t}$. During training, $\Delta W_{t}$ is stored separately from $W_{0}$ as two low-rank matrices with rank $r$: $\Delta W_{t}=B_{t}A_{t}$ with $B_{t}\in\mathbb{R}^{d\times r},A_{t}\in\mathbb{R}^{r\times k},W_{0}\in\mathbb{R}^{d\times k}$. $r$ is typically very small, often between $4$ and $32$. $\alpha$ is a constant that controls how much the LoRA adapters affect the behavior of the base weights $W_{0}$.
We hypothesize that LoRA’s low-rank updates restrict the model’s capacity to memorize precise details, which could have an effect similar to the noisy updates in DP. Therefore, LoRA has the potential to provide better privacy-utility trade-offs than FFT, similar to DP, while also being computationally more efficient.
Privacy-Utility trade-off: We investigate the effects of varying both the rank $r$ and scaling parameter $\alpha$ of LoRA. We use common rank values of $16$ and $32$ and $\alpha\in\{16,32,64,128\}$. While LoRA has been explored for privacy in conjunction with DP [51], there has been no prior work that specifically examines the privacy benefits of LoRA alone.
Figures 11(a)11(c) show the trade-off with rank $16$ and varying parameters of $\alpha$ for the CustomerSim dataset across Pythia, Gemma and Llama2 models. As seen in Figures 11(a) and 11(b), smaller-scale models (Pythia and Gemma) exhibit a better privacy-utility trade-off when the rank and $\alpha$ values are equal, compared to the other configurations. For the larger model, the privacy declines at later epochs, while utility is mostly retained. This trend is similar to the one observed for rank $32$ in Figures 11(d)11(f). We also analyze the trade-off for the SynBio dataset in Figures 12(a)12(c) for rank 16, and Figures 12(d)12(f) for rank 32, which make similar observations. However, due to the more unstructured nature of the SynBio dataset, there is a larger reduction in utility after certain epochs compared to the CustomerSim dataset.
Impact on benchmark datasets: We use the LoRA model with configuration $r=16,\alpha=16$ for this experiment. Figures 13(a)13(c) illustrate the LoRA fine-tuned model’s accuracy at each epoch for the three benchmarks. The LoRA-based fine-tuned Gemma model retains performance levels close to those of the pre-trained model.
Efficiency: During training, LoRA has slightly larger compute requirements for the forward pass than full fine-tuning, since additional FLOPs are required for the adapter matrices $\Delta W$, though they are much smaller than the full base weight $W_{0}$. However, during the backward pass, LoRA requires less compute, since no gradients need to be computed for the base weights $W_{0}$. We observe a relative FLOPs requirement of $C_{\text{LoRA}}/C_{\text{FFT}}\approx 0.65$. The original paper [20] reports a 25% speedup during training. LoRA has needs of less GPU memory than FFT, since no optimizer states and gradients need to be stored for the base weights $W_{0}$, which makes it possible to run it with larger batch-sizes and thus an overall increased training throughput.
(a) Pythia (CustomerSim)
(b) Gemma (CustomerSim)
(c) Llama2 (CustomerSim)
(d) Pythia (SynBio)
(e) Gemma (SynBio)
(f) Llama2 (SynBio)
Figure 14: Privacy-utility tradeoffs for different fine-tuning methods on the CustomerSim and SynBio datasets. Top row: On the CustomerSim dataset, DP achieves high privacy with a significant computational cost, full fine-tuning maximizes utility but offers less privacy, and LoRA provides a balanced trade-off with the lowest computational overhead. Bottom row: Similar trends are observed for the SynBio dataset.
Takeaway: We are one of the first works to explore the privacy benefits of parameter-efficient fine-tuning methods, particularly LoRA. We vary LoRA’s $r$ and $\alpha$ hyperparameters and observe that all configurations are able to achieve high utility, while especially lower $r$ and $\alpha$ values also preserve a high degree of privacy. For each model, the optimal privacy-utility trade-off value is achieved with $r=\alpha$. In addition, we observe that LoRA, after being fine-tuned on the CustomerSim dataset, did not lose much of its abilities on the benchmark datasets, maintaining results comparable to those of pre-trained model. Finally, LoRA is much more computationally and memory efficient than FFT and especially DP. Overall it provides the best trade-offs in terms of utility, privacy and efficiency and shows that privacy can be achieved without additional computational costs. The degree of measures along the trade-off, knowledge retention, and efficiency are:
Utility-privacy trade-offs: good
Retention of base performance: moderate
Efficiency: good
5 Comparison of Fine-tuning Methods
In this section, we compare all fine-tuning methods across three key aspects: privacy, utility, and efficiency. Recall that privacy is measured as training loss on sensitive tokens and utility is measured as test loss on non-sensitive tokens. For efficiency, we measure floating point operations (FLOPs) based on the number of operations incurred (e.g., matrix multiplication, addition, etc.) during training. In this section, for DP we use a noise ratio of $\sigma=0.1$ and for LoRA we use $r=16,\alpha=16$. Figures 14(a)-14(c) present the Pareto plots for CustomerSim, comparing the three fine-tuning strategies across three dimensions: privacy, utility, and efficiency.
Privacy: Regarding privacy, as previously noted, full fine-tuning shows poor privacy performance over extended training, while DP achieves the highest privacy levels. LoRA provides a similar degree of privacy compared to DP throughout most training epochs for small-scale models (Pythia and Gemma), but declines with the large-scale model (Llama2).
Utility: In terms of utility, full fine-tuning maintains relatively strong performance throughout most of its training (right on the x-axis). DP while yielding good utility in smaller models (Pythia and Gemma), performs poorly in the larger model (Llama2) and overall needs more epochs to reach higher utility levels. In contrast, LoRA consistently preserves higher utility across the entire training period.
Efficiency: The color bar in Figure 14 highlights the FLOPs intensity associated with each model across all fine-tuning strategies. DP requires the highest number of FLOPs (in red) due to the need for per-sample gradient computation, where each sample corresponds to a token within each training sequence. FFT demands a moderate number of FLOPs (in blue), proportional to the total number of parameters. Finally, LoRA requires the fewest FLOPs (in green), as especially during backpropagation, most operations only operate on the low-rank matrices.
(a) Benchmark SCIQ
(b) Benchmark MMLU
(c) Benchmark HellaSwag
Figure 15: LoRA demonstrates superior knowledge retention on benchmark datasets throughout the training regime on the CustomerSim dataset, while FFT and DP show fragility with sharp and gradual performance declines, respectively.
Benchmark performance: Figure 15 shows strong knowledge retention capabilities of LoRA on the three benchmarks after being fine-tuned on the CustomerSim dataset for 50 epochs. FFT and DP, on the other hand, decline sharply and gradually, respectively from the pretrained base model performance.
Assessing all the above aspects, we can observe the following:
1. Full fine-tuning achieves high utility initially, but starts diminishing after a few epochs, and also witnesses a significant drop in privacy. It has a relatively high computational cost. Additionally, the fully fine-tuned model’s performance diminishes significantly on benchmark datasets.
2. DP offers the strongest privacy protection and achieves a reasonable privacy-utility tradeoff in smaller models. However, this tradeoff deteriorates in larger models. DP incurs the highest computational cost as its per-sample noisy gradient updates significantly increase FLOPs and also memory requirements. Additionally, models fine-tuned with DP exhibit a gradual decline in their benchmark performance over the course of training.
3. LoRA, a parameter-efficient fine-tuning method, maintains high utility and achieves privacy levels comparable to DP in smaller models, though this advantage reduces in larger ones. Figure 14(c) shows that while LoRA preserves less privacy as training progresses, it is possible to select checkpoints that balance strong privacy with utility. This finding challenges the prevailing notion that privacy must come at the cost of high efficiency, demonstrating that LoRA can offer privacy benefits. Moreover, LoRA-tuned models retain performance on benchmark datasets close to that of pre-trained models throughout training. Figure 5 in Appendix 16 shows a comprehensive comparison across all the fine-tuning methods.
6 Concluding Discussion
In this paper, we examine the trade-offs among privacy, utility, and efficiency while fine-tuning an LLM. The traditional wisdom of achieving privacy comes at the cost of computational inefficiency using dedicated methods like DP. In contrast, we demonstrate that parameter efficient fine-tuning methods like LoRA, initially designed for efficiency, achieves privacy of sensitive data without any computational overhead. Simultaneously, LoRA retains the utility of general language understanding compared to DP, or even full-fine-tuning, realizing the superiority of LoRA in optimizing all three aspects. Towards our investigation, we establish the significance of redefining privacy and utility using a careful distinction between sensitive and non-sensitive counterparts of the fine-tuned data. Through case studies, we demonstrate how existing measures exaggerate privacy threats and undermine the utility of an LLM. Our paper calls for a joint venture of privacy and systems communities in achieving privacy-aware efficient fine-tuning of LLMs while retaining utility.
7 Ethics considerations
This study utilized publicly available datasets [44, 18], some of which included identifiable information such as personal details. However, third-party organizations pre-processed and validated the datasets to ensure that no real individuals’ data were present, thus mitigating potential privacy concerns.
This project received ethical clearance from the Ethical Review Board of the affiliated institution on October 21, 2024 (Approval No. 24-09-4), with no ethical concerns raised.
8 Open science
This work promotes transparency and reproducibility in research on privacy and utility in large language models (LLMs). To enable further investigation, we will release:
  • 1.
    Code and Framework: The implementation of our proposed privacy measurement framework, which distinguishes between sensitive and non-sensitive tokens, along with scripts for privacy leakage analysis and utility-efficiency evaluation.
  • 2.
    Datasets and Preprocessing information: Links to publicly available datasets used in this study, along with preprocessing scripts to ensure reproducibility. Sensitive data were excluded or anonymized to comply with ethical standards.
  • 3.
    Evaluation Pipeline: An open-source pipeline for assessing privacy leakage and the trade-offs between privacy, utility, and efficiency in LLMs.
These resources aim to support reproducibility and further research into privacy-aware, efficient LLM development.
Appendix
9 Hyperparameters
The following hyperparameters were used for fine-tuning our models:
Table 3: Hyperparameters used for fine-tuning methods
Hyperparameter Full Fine-Tuning (FFT) LoRA DP
Learning Rate 0.00025 0.00025 0.00005
Scheduler Linear Linear Linear
Warmup Steps 10 10 10
Clipping gradient - - ($1\times 10^{-2}$)
Table 4: Batch sizes for CustomerSim and SynBio
Dataset Model Full Fine-Tuning (FFT) LoRA DP
CustomerSim Pythia 16 32 8
Gemma 8 16 4
Llama2 4 8 2
SynBio Pythia 8 16 4
Gemma 4 8 2
Llama2 2 4 2
10 Validating GPT-4 Predictions
We analyzed the predictions shown in the heatmap in Figure 16 and observed that misclassified instances were frequently assigned to sections of a similar nature (e.g., NIH Explorer misclassified as PubMed Central). This overlap suggests that the nature of the misclassifications may not always indicate clear inaccuracies, making it difficult to definitively assess the accuracy of certain misclassified instances.
Figure 16: Misclassified GPT-4 instances were often placed in similar sections (e.g., NIH Explorer misclassified as PubMed Central), suggesting these misclassifications may not reflect clear inaccuracies, complicating definitive accuracy assessment.
11 Prompt for Annotating Privacy-Sensitive Information
We used the following prompt to obtain the privacy-sensitive information from GPT-4o for CustomerSim: {mdframed}[backgroundcolor=lavender, linewidth=0pt] “You are a private data identifying bot. You will be provided a conversation between a user and a customer service bot. You need to identify ALL instances of private data in the conversation. Private data includes any information that can be used to identify an individual, such as names, phone numbers, email addresses, and locations. It can also include any other sensitive information, such as credit card numbers, social security numbers, tracking data, health information etc.. Identify all of these use your discretion to identify any other information that can be used to identify an individual.”
12 Human Survey: Assessing Privacy-Sensitivity in Memorized Segments
We conducted a human annotation survey to assess if instances memorized by an LLM contain privacy-sensitive information, sampling 100 sequences divided into 5 questionnaires. Each was given to 40 Prolific workers. After an introductory explanation, participants evaluated each instance by (1) flagging privacy-sensitive information (ignoring routine commands unless uniquely identifiable) and (2) listing detected items if flagged. Figure 6 shows that most questions were labeled as non-sensitive.
13 Human Survey: Assessing the Quality of Privacy-Sensitive Annotations from Presidio and GPT-4
We evaluated privacy-sensitive annotations from Presidio and GPT-4 through two Prolific surveys, each with 10 samples for 40 participants. Binary Survey: For each sample, participants viewed side-by-side annotations from both tools (order randomized) and chose the more accurate one. Figure 17(a) shows GPT-4 was consistently preferred over Presidio for identifying privacy-sensitive information. Multiple-Choice Survey: Each of the 10 instances included 5 samples from each tool, displayed randomly. Participants had four options: (1) Accurate (all sensitive info correctly annotated), (2) Under-Annotated (some sensitive info missed), (3) Over-Annotated (non-sensitive info included), and (4) Mixed-Annotated (both missed and non-sensitive info annotated). Figure 17(b) shows GPT-4 rated as “Accurate” across datasets.
(a) Binary survey
(b) Multiple-choice survey
Figure 17: (a) GPT-4 was also preferred in SynBio by human annotators for identifying privacy-sensitive information. (b) GPT-4 also demonstrates in SynBio higher accuracy than Presidio for identifying privacy-sensitive information.
14 Why distinguish between sensitive and non-sensitive entities?
Figure 18 shows the plots observed in Figure 2 in log-scale for better distinction.
(a) Privacy measure in Pythia
(b) Utility measure in Pythia
(c) Privacy measure in Gemma
(d) Utility measure in Gemma
(e) Privacy measure in Llama2
(f) Utility measure in Llama2
Figure 18: Our measures offer a more precise assessment of privacy and utility when fine-tuning LLMs by distinguishing between sensitive and non-sensitive tokens, revealing higher privacy (higher loss) for sensitive tokens and better utility (lower loss) for non-sensitive tokens compared to traditional measures that overlook this sensitivity-based distinction.
15 Exploring the Tradeoffs between Privacy and Utility
We see a similar trade-off between utility and privacy when using Presidio annotations in Figure  19, as observed in Figure 7. For the CustomerSim dataset, we notice that privacy decreases at the beginning of training, while utility improves. However, as training progresses, both metrics start to worsen. In contrast, the SynBio dataset shows a decline in both utility and privacy, except for the Pythia model, which initially experiences an improvement in utility during the first few epochs.
(a) CustomerSim
(b) SynBio
Figure 19: Full fine-tuning on (a) CustomerSim and (b) SynBio using Presidio to annotate the privacy-sensitive information.
A similar pattern as in Figure 9 using GPT-4 annotations, has also emerged in figure 20 using Presidio annotations. We observe that in the CustomerSim dataset 20, DP maintains privacy effectively with minimal degradation. However, when it comes to utility, we notice that lower noise values lead to better utility, but also result in a decrease in privacy. The same holds true for the SynBio dataset 21, with the only exception being Gemma experiencing a decline in utility after a few iterations.
(a) Pythia
(b) Gemma
(c) Llama2
Figure 20: DP on CustomerSim dataset using Presidio tool for annotating privacy-sensitive information.
(a) Pythia
(b) Gemma
(c) Llama2
Figure 21: DP on SynBio dataset using Presidio tool for annotating privacy-sensitive information.
Comparing the results from Figures 22 and 23 using Presidio annotations with Figure 11, which used GPT-4 annotations, we can observe the same trend. This suggests that the findings are consistent across different annotation methods. Similarly, when we analyze the trade-off for the SynBio dataset in Figures  24 and 25, we see similar observations to those discussed in Figure 12, using GPT-4 annotations.
(a) Pythia
(b) Gemma
(c) Llama2
Figure 22: Lora with rank 16 on CustomerSim dataset using Presidio tool for annotating privacy-sensitive information.
(a) Pythia
(b) Gemma
(c) Llama2
Figure 23: Lora with rank 32 on CustomerSim dataset using Presidio tool for annotating privacy-sensitive information.
(a) Pythia
(b) Gemma
(c) Llama2
Figure 24: Lora with rank 16 on SynBio dataset using Presidio tool for annotating privacy-sensitive information.
(a) Pythia
(b) Gemma
(c) Llama2
Figure 25: Lora with rank 32 on SynBio dataset using Presidio tool for annotating privacy-sensitive information.
Analyzing the trade-offs between privacy, utility, and efficiency in Figures 26 and 27 for CustomerSim and SynBio using Presidio annotations reveals the same efficiency ranking observed in Figure 14 with GPT-4 annotations. LoRA is the most efficient method, while DP is the least efficient, requireing the highest number of FLOPs.
(a) Pythia
(b) Gemma
(c) Llama2
Figure 26: Full fine-tuning, DP and LoRA on CustomerSim with Presidio annotations for privacy-sensitive information.
(a) Pythia
(b) Gemma
(c) Llama2
Figure 27: Full fine-tuning, DP and LoRA on SynBio with Presidio annotations for privacy-sensitive information.
16 Comparison of fine-tuning methods
Table 5: Comparison of all fine-tuning methods across privacy, utility, efficiency, and benchmark performance. LoRA is seen to outperform FFT and DP over all dimensions.
Utility Privacy Efficiency Benchmark
FFT Good Poor Moderate Poor
DP Moderate Good Poor Poor
LoRA Good Good Good Good